OpenStudy (anonymous):

How do you solve this??? The solution set of the equation |x-1|=3 is

OpenStudy (anonymous):

x=4 me bad

OpenStudy (anonymous):

how do you solve it?

OpenStudy (anonymous):

the same as any expression What minus 1 will give you a numbr 3 spots away from 0 so the answer is four

OpenStudy (anonymous):

since there is an absolute value there are two answers for this... solve once for x-1 = 3 and once for x-1 = -3

OpenStudy (anonymous):

|x-1|=3 x-1=3 and x-1=-3 x=4 and x=-2